I need to solve this equation by using the Substitution method
x + y = 7
2x + 2y = 14

Answer by Ed Parker(21)

If you try to solve it by the substitution method:

Solve the first equation for y

x + y = 7
    y = 7 - x

Substitute y = 7 in the second equation:

      2x + 2y = 14
2x + 2(7 - x) = 14
 2x + 14 - 2x = 14
           14 = 14

The x's canceled out and left a true numerical equation 14 = 14.  

So it is a dependent system. It has infinitely many solutions.

For instance 

1.  x=3, y=4, the ordered pair (x,y) = (3,4) is a solution:

Checking in the first       Checking in the second
equation:                   equation:

 x + y = 7                      2x + 2y = 14
 3 + 4 = 7                  2(3) + 2(4) = 14
     7 = 7                        6 + 8 = 14
                                     14 = 14

2.  x=2, y=5, the ordered pair (x,y) = (2,5) is a solution:

Checking in the first       Checking in the second
equation:                   equation:

 x + y = 7                      2x + 2y = 14
 2 + 5 = 7                  2(2) + 2(5) = 14
     7 = 7                       4 + 10 = 14
                                     14 = 14 

3.  x=-3, y=10, the ordered pair (x,y) = (-3,10) is a solution:

Checking in the first       Checking in the second
equation:                   equation:

 x +  y = 7                      2x + 2y = 14
-3 + 10 = 7                2(-3) + 2(10) = 14
      7 = 7                      -6 + 20 = 14
                                      14 = 14

I could show you thousands and thousands of solutions because the
two equations have the same line as a graph, and so they have ALL
points in common.

Answer:  {(x,y)|y = 7-x}  dependent system  

Edwin